# There and Back Again: On the relation between Noise and Image Inversions in Diffusion Models

### v4

@@ Appendix E Approximation error in DDIM Inversion
Appendix G Conditioning Diffusion Models
For a thorough analysis, we employ both unconditional (ADM-32, ADM-64, ADM-256) and conditional (DiT, IF, SDXL) diffusion models. For conditioning, we take prompts from the Recap-COCO-30K dataset (Li et al., 2025) for IF and Stable Diffusion XL and ImageNet class names for DiT. However, as noted by Mokady et al. (2023), Classifier-Free Guidance introduces additional errors to the DDIM inversion, so to focus solely on the inversion approximation error, we disable CFG by setting the guidance scale to $w=1$.
Appendix H Most probable triangles
H.1 Methodology
H.2 Results
Model $T$ $\angle{\mathbf{x_{0}}}$ $\angle{\mathbf{x_{T}}}$ $\angle{\mathbf{\hat{x}_{T}}}$
ADM
$32\times 32$
$10$ $44$ $16$ $120$
$100$ $29$ $28$ $123$
$1000$ 2020 $45$ $115$
ADM
$64\times 64$
$10$ $30$ $31$ $119$
$100$ $11$ $60$ $109$
$1000$ 66 $79$ $95$
ADM
$256\times 256$
$10$ $24$ $50$ $106$
$100$ $24$ $73$ $83$
$1000$ $23$ $73$ $84$
LDM
$256\times 256$
$10$ $23$ $53$ $104$
$100$ 22 $76$ $102$
$1000$ 11 $83$ $96$
DiT
$256\times 256$
$10$ $27$ $47$ $106$
$100$ 44 $66$ $110$
$1000$ 11 $80$ $99$
(a) ADM-32 (CIFAR-10)
(b) ADM-256 (ImageNet)
(c) LDM (CelebA)
H.3 Example histograms
Appendix I Plain surface thresholding
During the experiments, we set $\tau=0.025$.
$\mathcal{M}(h,w)=\prod_{c}\mathcal{M}_{c}(h,w).$ (12)
Appendix J Calculating inversion error
Further, for the images in the dataset, we obtain binary masks $\mathcal{M}_{p}$ and $\mathcal{M}_{n}$ indicating, respectively, plain ($p$) and non-plain ($n$) pixels in the image, according to the procedure described in Appendix I. To ensure that the level of noise that DM predicts in each step does not bias the results, we divide the absolute errors in each step by $l_{1}$-norm of model outputs, calculated separately for each diffusion step. For plain ($p$) pixels, it can be described as
and adequately for non-plain ($n$) pixels as
Pixel area Diffusion steps Model
ADM-64 DiT
Plain $1,2\dots 50$ $100\%$ $15.11$ $5.67$
Non-plain $12.43$ $3.16$
Plain $1,2\dots 5$ $10\%$ $9.23$ $3.41$
Non-plain $6.80$ $1.42$
Plain $6,7\dots 50$ $90\%$ $5.87$ $2.23$
Non-plain $5.63$ $1.74$
Appendix K Pseudocode for DDIM Inversion with Forward Diffusion
We present in Algorithm 1 the pseudocode for the proposed solution to the decorrelation of latent encodings by replacing the first $t^{\prime}$ inversion steps with a forward diffusion process.
2:decorrelated latent encoding $\mathbf{\hat{x}_{T}}$}
3:sample $\tilde{\epsilon}\sim\mathcal{N}(0,\mathcal{I})$
[…]
5:for $t\leftarrow f+1,\dots,T$ do
7:end for
8:return $\mathbf{\hat{x}_{T}}$}
Appendix L Details on diversity and alignment of edition
Inversion steps replaced Diversity against $I_{S}$ CLIP Alignment
Noise $\mathbf{x_{T}}$} $0.709$ $0.591$ $0.228$ $0.174$ $0.465$ $0.681$
DDIM Latent $\mathbf{\hat{x}_{T}}$} $0.677$ $0.564$ $0.261$ $0.223$ $0.480$ $0.662$
11 ($2\%$) $0.696$ $0.580$ $0.239$ $0.187$ $0.470$ $0.676$
$1\dots 2$ ($4\%$) $0.701$ $0.587$ $0.227$ $0.179$ $0.469$ $0.677$
$1\dots 3$ ($6\%$) $0.705$ $0.594$ $0.216$ $0.173$ $0.468$ $0.678$
$1\dots 5$ ($10\%$) $0.711$ $0.605$ $0.198$ $0.165$ $0.466$ $0.679$
$1\dots 50$ ($100\%$) $0.805$ $0.801$ $0.013$ $0.085$ $0.465$ $0.680$
(a) Diffusion Transformer
Inversion steps replaced Diversity against $I_{S}$ CLIP Alignment
Noise ($100\%$) $0.665$ $0.380$ $0.339$ $0.344$ $0.273$ $0.649$
DDIM Latent $\mathbf{\hat{x}_{T}}$} $0.609$ $0.328$ $0.406$ $0.416$ $0.353$ $0.614$
11 ($2\%$) $0.666$ $0.376$ $0.359$ $0.335$ $0.281$ $0.646$
$1\dots 2$ ($4\%$) $0.660$ $0.369$ $0.359$ $0.351$ $0.287$ $0.649$
$1\dots 3$ ($6\%$) $0.661$ $0.370$ $0.353$ $0.349$ $0.285$ $0.650$
$1\dots 5$ ($10\%$) $0.662$ $0.372$ $0.341$ $0.346$ $0.282$ $0.649$
$1\dots 50$ ($100\%$) $0.733$ $0.521$ $0.004$ $0.272$ $0.274$ $0.649$
(b) Deepfloyd IF
Appendix M Stochastic editing of real images
In this work, we propose the solution for decorrelating latent encodings resulting from DDIM inversion by replacing its first steps with the forward diffusion. As presented in Algorithm 1, the forward diffusion process involves sampling random Gaussian noise $\tilde{\epsilon}\sim\mathcal{N}(0,\mathcal{I})$ and interpolating it with input image. Due to the fact that, when we replace a small fraction of steps ($2-4\%$), the change in image reconstruction error is insignificant, the use of different noises $\tilde{\epsilon}$ (in practice, sampled with different seeds) allows stochastic image editing, i.e. generating different manipulations of the input image, a feature not naturally available with DDIM inversion. In Fig. 13, we present examples of editing real images from the ImageNet-R-TI2I dataset (which we annotate using GPT-4o) with the IF model, showing various semantically correct modifications of the same image.
As preserving original image structure during editing is stated as a more difficult task for real images than the one naturally generated by diffusion model, we follow Hertz et al. (2022) by, first, denoising latent encodings with source prompt (the one used during inversion), and, after $6\%$ of the steps, using target prompt as conditioning. The examples presented in Fig. 13 indicate, that our fix (1) enables stochastic editing of images and (2) enables image manipulations in plain image regions, contrary to editing with naïve DDIM latents.
Figure 13: Replacing first DDIM inversion steps with forward diffusion enables stochastic image editing, resulting in multiple semantically correct manipulations of same input image. Contrary to DDIM Inversion, editing with latents produced by the solution introduced in this work, enables image manipulations in uniform input image areas.
Appendix N Noise-to-image mapping
We showcase an additional study showing the differences that occur between noise and latent encodings, from the perspective of their mapping to the images. Several works investigate interesting properties between the initial random noise and generations that result from the training objective of DDPMs and score-based models. Kadkhodaie et al. (2024) show that due to inductive biases of denoising models, different DDPMs trained on similar datasets converge to almost identical solutions. This idea is further explored by Zhang et al. (2024a), observing that even models with different architectures converge to the same score function and, hence, the same noise-to-imape mapping. Khrulkov & Oseledets (2022) show that diffusion models’ encoder map coincides with the optimal transport (OT) map when modeling simple distributions. However, other works (Kim & Milman, 2011; Lavenant & Santambrogio, 2022) contradict this finding.
N.1 Smallest $l_{2}$ mapping
[…]
T ADM-32 ADM-64 ADM-256 LDM DiT
$4000$ $99.1_{\pm 1.0}$ $46.4_{\pm 3.0}$ $99.5_{\pm 0.3}$ $43.3_{\pm 6.7}$ - - - - - -
T ADM-32 ADM-64 ADM-256 LDM DiT
$4000$ $2.8_{\pm 2.2}$ $41.9_{\pm 3.0}$ $1.9_{\pm 1.4}$ $43.5_{\pm 6.5}$ - - - - - -
[…]
N.2 Asymmetry of Noise-to-Image mapping
Conclusion.
Appendix O On Noise-Image-Latent relations during diffusion training
To further explore the relationships that exist between noises, generations, and latents, we study how the relationships between them change with the training of the diffusion model. We train two diffusion models from scratch and follow the setup from Nichol & Dhariwal (2021) for two unconditional ADMs for the ImageNet ($64\times 64$) and CIFAR-10 ($32\times 32$) datasets. The CIFAR-10 model is trained for 700700K steps, while the ImageNet model – for 1.51.5M steps, both with a batch size of $128$. Models employ a cosine scheduler with 44K diffusion steps.
O.1 Spatial relations of noise and latents over training time
O.2 Image-to-noise distance mapping over training time
We analyze the image-noise mapping with $l_{2}$-distance from Appendix N over diffusion model training time. We sample $N=2000$ Gaussian noises and generate images from them using ADM models with $T=100$ diffusion steps, calculating the accuracy of assigning images to corresponding noises (and vice versa) using the smallest $l_{2}$-distance. In Fig. 17, we can observe, for both models, that the distance between noises and their corresponding generations accurately defines the assignment of initial noises given the generated samples (${\mathbf{x_{0}}}\to{\mathbf{x_{T}}}$) from the beginning of the training till the end. At the same time, the accurate reverse assignment (${\mathbf{x_{T}}}\to{\mathbf{x_{0}}}$) can only be observed at the beginning of the training when the trained model is not yet capable of generating properly formed images. Already in the beginning phase of model training, the quality of noise to image mapping rapidly drops and does not change until the end of training.
(a) ADM-64
(b) ADM-32
O.3 Image alignment over training time
In Fig. 19, we visualize how the diffusion model learns the low-frequency features of the image already at the beginning of the fine-tuning when comparing generations from the next training steps against the generations after finished fine-tuning for the ADM-32 model trained on the CIFAR-10 dataset. In  Fig. 20 (ADM-64) and Fig. 21 (ADM-32), we show additional examples illustrating how generations evolve over training for the same Gaussian noise $\mathbf{x_{T}}$} using a DDIM sampler. Initially, low-frequency features emerge and remain relatively stable, while continued fine-tuning improves generation quality by refining only the high-frequency details.
Figure 19: Similarity of the generations sampled from the same random noise at different stages of diffusion model’s training to the final outputs for ADM-32 (CIFAR-10). We plot CKA, SVCCA, SSIM, and DINO image alignment metrics and show that the diffusion model already learns the mapping between Gaussian noise and generations at the beginning of the training.
Appendix P Additional results
In this section, we present qualitative results of replacing first DDIM Inversion steps with forward diffusion.
P.1 Failure cases
As described in Section 4, while replacing a higher number of inversion steps with forward diffusion enables better editability of latent encodings (through decorrelating them), it increases the image reconstruction error. In Fig. 22, we show that replacing $10\%$ or $20\%$ of first steps can introduce significant changes to images.
Figure 22: Replacing first DDIM inversion steps with forward diffusion increases latents editability at cost of a higher reconstruction error. By replacing from $2\%$ up to $4\%$ of steps we obtain reasonable image reconstructions while removing correlations from latents.
P.2 Interpolation
In Section 4, we presented that interpolating DDIM latents with SLERP (Shoemake, 1985) leads to decrease in image quality and diversity when compared to Gaussian noise. In Fig. 23, we qualitatively compare our fix for removing correlations in latent encodings with naïve DDIM inversion in the task of image interpolation.
Figure 23: Qualitative comparison of images generated from interpolated latents produced with DDIM Inversion and our fix. Contrary to naïve DDIM inversion, proposed solution enables generating high quality objects with pixel-diverse background.

### v5

@@ Conclusion.
N.2 Percentage of inversion steps replaced
The fix to DDIM Inversion algorithm, proposed in this work, namely replacing neural network predictions with random Gaussian Noise, implies the trade-off between preserving the original image information and improving the latents’ editability. In this section, we present how the number of inversion steps substituted with forward step, impacts the image reconstruction error (MAE, LPIPS, and SSIM metrics) and latent editability (correlations and KL Divergence from $\mathcal{N}(0;\mathcal{I})$) for: IF (Table 15), DiT (Table 16), and SDXL (Table 17) models. We observe that replacing only the first $4\%$ of inversion steps with forward diffusion improves latent normality to the level of Gaussian noise ($100\%$ of steps), while increases reconstruction error only slightly. In Fig. 22, we show some failure cases when replacing $10\%$ or $20\%$ of the first steps can introduce significant changes to images.
# Steps Replaced Image Reconstruction Latent Normality
(Percentage) MAE $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ Correlation $\downarrow$ KL Div. $\times 10^{2}$ $\downarrow$
0 (0%) 0.073 0.030 0.878 0.643 60.449
1 (2%) 0.069 0.037 0.854 0.057 0.934
2 (4%) 0.071 0.038 0.845 0.050 0.352
3 (6%) 0.074 0.040 0.830 0.050 0.346
4 (8%) 0.078 0.043 0.813 0.049 0.341
5 (10%) 0.082 0.047 0.796 0.049 0.338
10 (20%) 0.099 0.066 0.713 0.049 0.344
20 (40%) 0.131 0.113 0.556 0.049 0.360
30 (60%) 0.169 0.179 0.394 0.049 0.367
40 (80%) 0.233 0.279 0.204 0.049 0.370
50 (100%) 0.487 0.437 0.009 0.049 0.374
Table 15: Impact of percentage of inversion steps replaced with forward diffusion on reconstruction quality and latent normality for DeepFloyd IF.
# Steps Replaced Image Reconstruction Latent Normality
(Percentage) MAE $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ Correlation $\downarrow$ KL Div. $\times 10^{2}$ $\downarrow$
0 (0%) 0.052 0.063 0.839 0.159 1.118
1 (2%) 0.070 0.097 0.741 0.038 0.011
2 (4%) 0.085 0.125 0.658 0.036 0.010
3 (6%) 0.097 0.151 0.594 0.036 0.023
4 (8%) 0.107 0.173 0.544 0.037 0.036
5 (10%) 0.116 0.195 0.505 0.037 0.041
10 (20%) 0.154 0.280 0.375 0.038 0.040
20 (40%) 0.231 0.437 0.235 0.037 0.020
30 (60%) 0.353 0.595 0.145 0.037 0.017
40 (80%) 0.521 0.710 0.071 0.037 0.017
50 (100%) 0.628 0.750 0.029 0.037 0.018
Table 16: Impact of percentage of inversion steps replaced with forward diffusion on reconstruction quality and latent normality for Diffusion Transformer (DiT).
# Steps Replaced Image Reconstruction Latent Normality
(Percentage) MAE $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ Correlation $\downarrow$ KL Div. $\times 10^{2}$ $\downarrow$
0 (0%) 0.027 0.099 0.814 0.166 0.800
1 (2%) 0.029 0.106 0.790 0.151 0.600
2 (4%) 0.035 0.137 0.716 0.120 0.300
3 (6%) 0.038 0.155 0.685 0.117 0.300
4 (8%) 0.041 0.171 0.663 0.116 0.300
5 (10%) 0.043 0.183 0.646 0.116 0.200
10 (20%) 0.051 0.230 0.588 0.115 0.200
20 (40%) 0.065 0.313 0.515 0.115 0.200
30 (60%) 0.082 0.389 0.460 0.115 0.200
40 (80%) 0.112 0.464 0.464 0.115 0.200
50 (100%) 0.159 0.541 0.541 0.115 0.200
Table 17: Impact of percentage of inversion steps replaced with forward diffusion on reconstruction quality and latent normality for Stable Diffusion XL.
Figure 22: Replacing first DDIM inversion steps with forward diffusion increases latents editability at the cost of a higher reconstruction error. By replacing from $2\%$ up to $4\%$ of steps, we obtain reasonable image reconstructions while removing correlations from latents.
N.3 Guidance scale
During experiments, we fix the guidance scale to $w=1$ to ensure that our analysis focuses solely on the DDIM approximation error (Eq. 3). As described in Ju et al. (2024), prior works typically employ guidance scale between $1.0$ (the most common choice) and $3.0$ during inversion, as using $w>3$ often results in drastically worse image reconstructions.
In this section, we evaluate how the proposed fix improves upon Naïve DDIM Inversion when a higher guidance scale $w\in\{1,2,3,4,5\}$ is applied both during inversion and reconstruction. In Table 18, we present result of this experiment with Stable Diffusion XL (Podell et al., 2024). For scenarios with a higher guidance scale, our simple fix, similarly to $w=1$, reduces correlations and improves editability when comparing with Naive DDIM Inversion. Additionally, we observe that, when guidance is applied, our fix improves image reconstruction error (measured with LPIPS). We hypothesize that amplifying the inversion error in Naïve DDIM with a higher guidance scale leads to latents that are useless for image reconstruction. In such a case, replacing the first steps with random Noise may lead to more preferable reconstructions. In Fig. 23 and Fig. 24, we present qualitative comparison in image reconstruction between Naïve DDIM Inversion and our approach.
Guidance Method LPIPS $\downarrow$ Latent CLIP Alignment
Scale $w$ Corr. $\downarrow$ (Edit Prompt) $\uparrow$
1.0 DDIM Inv. 0.100 0.166 0.695
w/ ours (4%) 0.137 0.120 0.722
$\Delta$ +0.037 -0.046 +0.027
2.0 DDIM Inv. 0.199 0.170 0.779
w/ ours (4%) 0.179 0.120 0.807
$\Delta$ -0.020 -0.050 +0.028
3.0 DDIM Inv. 0.390 0.171 0.764
w/ ours (4%) 0.267 0.121 0.815
$\Delta$ -0.123 -0.050 +0.051
4.0 DDIM Inv. 0.525 0.172 0.725
w/ ours (4%) 0.372 0.121 0.800
$\Delta$ -0.153 -0.051 +0.075
5.0 DDIM Inv. 0.582 0.174 0.687
w/ ours (4%) 0.452 0.121 0.770
$\Delta$ -0.130 -0.053 +0.083
Table 18: Performance in image reconstruction (LPIPS), inverted latent normality (correlations), and text alignment to edit prompt for different values of guidance scale. We show that our forward step replacement (4%) improves DDIM Inversion algorithm.
Figure 23: Examples of image reconstruction with Naïve DDIM Inversion and DDIM Inversion incorporating our fix (forward $4\%$) for guidance scale $w=1$. Examples generated with Stable Diffusion XL.
Figure 24: Comparison of image reconstruction examples with Naïve DDIM Inversion and DDIM Inversion incorporating our fix (forward $4\%$) across different values of guidance scale $w\in\{2,3,4,5\}$. Examples generated with Stable Diffusion XL.
Appendix O Qualitative examples
O.1 Image interpolation
In Section 4, we presented that interpolating DDIM latents with SLERP (Shoemake, 1985) leads to a decrease in image quality and diversity when compared to Gaussian noise. In Fig. 25, we qualitatively compare our fix for removing correlations in latent encodings with naïve DDIM inversion in the task of image interpolation.
Figure 25: Qualitative comparison of images generated from interpolated latents produced with DDIM Inversion and our fix. Contrary to naïve DDIM inversion, the proposed solution enables generating high-quality objects with pixel-diverse backgrounds.
O.2 Reconstructions of real images
In Fig. 26, we present a qualitative comparison for reconstructions of real images from the StyleDrop (Sohn et al., 2023) dataset. We observe that DDIM Inversion with our fix sometimes provides imperfect image reconstructions. However, those failures are also observable with vanilla DDIM Inversion, indicating that they stem from DDIM approximation error itself, not from our replacement.
Figure 26: Examples of reconstructions of real images from the StyleDrop (Sohn et al., 2023) dataset with Naive DDIM Inversion and DDIM with our fix (forward diffusion in $4\%$ of steps). Inversion process is run with $T=50$ steps and guidance scale $w=1.0$. Reconstructions generated with Stable Diffusion XL.
O.3 Stochastic image editing
In this work, we propose a solution for decorrelating latent encodings resulting from DDIM inversion by replacing its first steps with the forward diffusion. As presented in Algorithm 1, the forward diffusion process involves sampling random Gaussian noise $\tilde{\epsilon}\sim\mathcal{N}(0,\mathcal{I})$ and interpolating it with the input image. Due to the fact that, when we replace a small fraction of steps ($2-4\%$), the change in image reconstruction error is insignificant, the use of different noises $\tilde{\epsilon}$ (in practice, sampled with different seeds) allows stochastic image editing, i.e. generating different manipulations of the input image, a feature not naturally available with DDIM inversion. In Fig. 27, we present examples of editing real images from the ImageNet-R-TI2I dataset (which we annotate using GPT-4o) with the IF model, showing various semantically correct modifications of the same image.
As preserving original image structure during editing is stated as a more difficult task for real images than the one naturally generated by the diffusion model, we follow Hertz et al. (2022) by, first, denoising latent encodings with source prompt (the one used during inversion), and, after $6\%$ of the steps, using target prompt as conditioning. The examples presented in Fig. 27 indicate that our fix (1) enables stochastic editing of images and (2) enables image manipulations in plain image regions, contrary to editing with naïve DDIM latents.
Figure 27: Replacing first DDIM inversion steps with forward diffusion enables stochastic image editing, resulting in multiple semantically correct manipulations of the same input image. Contrary to DDIM Inversion, editing with latents produced by the solution introduced in this work enables image manipulations in uniform input image areas.
O.4 Real image editing with MasaCtrl
In this section, we present examples for editing real images from the PIEBench dataset (Ju et al., 2024) when our inversion method is combined with the MasaCtrl (Cao et al., 2023) editing engine. In Figs. 28, 29 and 30, we qualitatively compare with Naïve DDIM Inversion across several editing tasks: object replacement, attribute editing, and object removal. We present that replacing first $4\%$ of inversion steps with forward diffusion leads to more successful edits in prompt adherence, while not observing degradation in consistency to input images.
Figure 28: Object replacement on real images with MasaCtrl. Comparison for Naïve DDIM Inversion and DDIM with our fix (forward diffusion in $4\%$ of steps). Model: Stable Diffusion 1.4.
Figure 29: Attribute editing on real images with MasaCtrl. Comparison for Naïve DDIM Inversion and DDIM with our fix (forward diffusion in $4\%$ of steps). Model: Stable Diffusion 1.4.
Figure 30: Object removal on real images with MasaCtrl. Qualitative comparison for Naïve DDIM Inversion and DDIM with our fix (forward diffusion in $4\%$ of steps). Examples generated with Stable Diffusion 1.4.
O.5 Style transfer with StyleAligned
In Figs. 31 and 32, we present a qualitative comparison of Naïve DDIM Inversion and our approach when combined with StyleAligned (Hertz et al., 2024) for the task of Style Transfer. Examples have been generated with Stable Diffusion XL using the same hyperparameters for both settings on the StyleDrop dataset (Sohn et al., 2023). We observe that replacing the first steps of DDIM Inversion with forward diffusion enables better prompt-adherence for generations.
Figure 31: Examples of style transfers from real images from the StyleDrop (Sohn et al., 2023) dataset. Comparison for Naive DDIM Inversion and DDIM with our fix ($4\%$ of steps replaced).
Figure 32: Examples of style transfers from real images from the StyleDrop (Sohn et al., 2023) dataset. Comparison for Naive DDIM Inversion and DDIM with our fix (forward diffusion in $4\%$ of steps). Examples generated with Stable Diffusion XL.
Appendix P Latent correlations in Flow Matching models
In this section, we analyze if, similarly to latents produced with DDIM inversion in Diffusion Models, inversion procedure with Flow Matching models leads to correlations.
The inversion procedure can be incorporated into Flow Matching (FM) models (Lipman et al., 2023; Liu et al., 2023), e.g., for image editing (Avrahami et al., 2025; Kulikov et al., 2025; Rout et al., 2025). The generative process of FMs is defined as an ordinary differential equation (ODE) over time $t\in[0,1]$ with time-dependent velocity field $V$:
$dz_{t}=V(z_{t},t)dt.$ (16)
Commonly, this ODE, given an initial condition $z_{1}\sim\mathcal{N}(0;\mathcal{I})$, is solved numerically with Euler method, leading to iterative sampling process $t\in\{T,T-1,\dots,1\}$, defined as
$z_{t-1}=z_{t}+(\sigma_{t-1}-\sigma_{t})\cdot\nu_{\theta}(z_{t},t),$ (17)
with $\nu_{\theta}$ being a neural network parametrizing the continuous velocity field leading to clean images $z_{0}$ and $\sigma_{t}$ being a noise schedule.
The inverse step, as described in Avrahami et al. (2025), can be expressed as
$z_{t}=z_{t-1}+(\sigma_{t}-\sigma_{t-1})\cdot\nu_{\theta}(\mathbf{z_{t-1}},t),$ (18)
with an assumption that locally $\nu_{\theta}(z_{t},t)\approx\nu_{\theta}(z_{t-1},t)$. We refer to this formulation as ODE Inversion.
As the approximation relies on a similar assumption as in the case of DDIM (Eq. 3), we analyze if the ODE Inversion, similarly, induces correlation patterns in output latents. In Table 19, we report image reconstruction error, editing textual alignment (CLIP Similarity to edit prompt and Directional Similarity (Gal et al., 2022)), and metrics validating the latents’ normality. We employ FLUX.1 (Labs, 2024) model with $T=50$ inversion and sampling steps. We present that the latents resulting from the ODE Inversion algorithm, similarly to the case of DDIM latents, exhibit correlations and visible deviation from the Gaussian distribution. Importantly, these deviations, when compared to using original noise, lead to a significant decrease in prompt alignment when starting the generation process with an editing prompt. Additionally, in Table 19, we compare original noise and ODE Inversion latent diversity for plain and non-plain input image pixel regions. Although not as significantly visible as in DDIM latents, ODE Inversion outputs as well tend to be more erroneous for plain image pixels and less diverse in those areas.
Finally, in Fig. 33, we present qualitative examples for image reconstructions and latent correlation when ODE Inversion is performed. As visible, after decoding with FLUX’s decoder, ODE Inversion latents exhibit correlations in locations that represent smooth pixel areas of images. Additionally, we plot the absolute error between original Gaussian Noise and ODE Inversion latents after applying PCA for dimensional reduction (as FLUX operates in $16$-channel latent space).
Metric Gaussian Noise ODE Inv. Latent
Image MAE $\downarrow$ 0.00 0.05
Reconstr. LPIPS $\downarrow$ 0.00 0.16
CLIP Text Edit prompt $\uparrow$ 81.49 56.40
Alignment Directional Sim. $\uparrow$ 87.94 55.32
Normality Correlation $\downarrow$ 0.14 0.27
KL Div. $\times 10^{-2}\downarrow$ 0.20 3.80
Noise Plain pixels 0.00 0.31
Error Non-plain pixels 0.00 0.26
Variance Plain pixels 0.98 0.94
Non-plain pixels 1.01 1.03
Table 19: Comparison between original Gaussian Noise and latents resulting from ODE Inversion process with FLUX.1 model. We show that ODE inversions are more correlated than Gaussian and significantly deviate from normal distribution. This leads to worse text alignment during editing.
Figure 33: ODE Inversion in Flow-Matching models, similarly as DDIM Inversion in Diffusion models, produces latent encodings with correlations. Reconstructions performed with FLUX.